OpenStudy (anonymous):

1 2/5 plus 5/8 equals what? i dont need to see the work my husband just wants to know he is out of town and trying to add something and I suck at math

OpenStudy (pfenn1):

\[1 \frac25+\frac 58=\frac75+\frac58=\frac{56+25}{(8)(5)}=\frac{81}{40}=2 \frac{1}{40}\]